# Can Generalist Vision Language Models (VLMs) Rival Specialist Medical VLMs? Benchmarking and Strategic Insights

### v2

@@ Abstract
Background: Vision–Language Models (VLMs) have shown promise in automating image diagnosis and interpretation in clinical settings. However, developing specialist medical VLMs requires substantial computational resources and carefully curated datasets, and it remains unclear under which conditions generalist and specialist medical VLMs each perform best.
Methods: This paper introduces , the first unified benchmark and evaluation framework designed for direct, fair comparisons of paired generalist and specialist medical VLMs from the same model families. Our benchmark is comprehensive, assessing 18 models across radiology, pathology, dermatology, and ophthalmology on 144 diagnostic and 72 visual-question answering (VQA) tasks. focusing on assessing both in-domain (ID) and out-of-domain (OOD) performance, with off-the-shelf and parameter-efficient fine-tuning (e.g., linear probing, LoRA). Diagnostic classification tasks were evaluated using AUROC, while VQA tasks were assessed with BLEU-1, ROUGE-L, Exact Match, F1 Score, and GPT-based semantic scoring, covering both open- and closed-ended formats.
Results: As expected, off-the-shelf specialist medical VLMs generally outperformed generalist VLMs on ID tasks given their pretraining. However, with lightweight fine-tuning, generalist VLMs achieved superior performance in most of ID task evaluations and demonstrated better generalization on OOD tasks in the majority of comparisons. Fine-tuning required only updating small amounts of parameters associated with full medical pretraining. In contrast, fine-tuned specialist medical VLMs do not show comparable performance even with the same level of fine-tuning. is fully open-sourced at https://github.com/ubc-tea/MedVLMBench.
Conclusions: This study highlights the complementary strengths of specialist medical and generalist VLMs. Specialists remain valuable in modality-aligned use cases, but we find that efficiently fine-tuned generalist VLMs can achieve comparable or even superior performance in most tasks, particularly when transferring to unseen or rare OOD medical modalities. These results suggest that generalist VLMs, rather than being constrained by their lack of specialist medical pretraining, may offer a scalable and cost-effective pathway for advancing clinical AI development.
Can Generalist Vision Language Models (VLMs) Rival Specialist Medical VLMs?
Benchmarking and Strategic Insights
1 Introduction
@@ 3 Results
Together, these two views allow us to address all our research questions. The complete raw data underlying these figures are included in Appendix D, with Tables 5–7 reporting diagnostic results and Tables 8–27 reporting VQA results.
@@ Appendix A Vision-Language Models
| LLaVA family | ||||||
| LLaVA-1.5 [liu2023visual] | Generalist | CLIP ViT-L/336 | Vicuna 7B | MLP projector; 1.2 M visual instructions (GPT-4, academic VQA) | ✓ | 7.38B |
| LLaVA-Med [li2023llava] | Specialist Medical | CLIP ViT-L/336 | Mistral-7B | Two-stage curriculum on 15 M PMC figure–caption pairs + GPT-4 self-instruct | ✓ | 7.90B |
| Qwen-VL family | ||||||
| Qwen-2-VL [wang2024qwen2] | Generalist | Dynamic-res ViT | Qwen-2 LLM | Web-scale image-text pairs, videos; SFT with dialogue, grounding, and agent data | ✓ | 8.29B |
| Qwen-2.5-VL [bai2025qwen2] | Generalist | Dynamic-res ViT | Qwen-2.5 LLM | Web-scale data; Multimodal Pre-Training, Long-Context Pre-Training, SFT, tool-use alignment | ✓ | 8.29B |
| Lingshu [xu2025lingshu] | Specialist Medical | Qwen-VL ViT | Qwen-2.5 (base) | Shallow align deep align med-instruction tune Reinforcement learning with verifiable rewards | ✓ | 8.29B |

### v3

@@ Abstract
Background Vision–language models (VLMs) are increasingly used to support automated medical image interpretation, yet the relative strengths of generalist versus specialist medical VLMs remain uncertain. Specialist models benefit from domain-specific pretraining but demand substantial computational resources, while generalist VLMs are widely available and easily adapted for clinical use. A systematic comparison of their performance across medical imaging tasks has been lacking.
Objective This study introduces MedVLMBench, a unified benchmark designed to evaluate generalist and specialist medical VLMs, and to examine how lightweight fine-tuning affects performance across in-distribution (ID) and out-of-distribution (OOD) diagnostic and visual question answering (VQA) tasks.
Methods and Analysis We assessed 18 VLMs, paired generalist and specialist variants across contrastive and generative model families, on ten datasets spanning radiology, pathology, dermatology, and ophthalmology. The benchmark comprises 144 diagnostic classification tasks and 72 VQA tasks. Off-the-shelf and parameter-efficient fine-tuned models (e.g., linear probing, LoRA) were evaluated using AUROC, BLEU-1, ROUGE-L, exact match, F1 score, and a GPT-based semantic scoring metric. Performance gaps were quantified to address three research questions concerning specialist advantage, the effect of lightweight adaptation, and generalization to OOD modalities. We open-source our code at https://github.com/ubc-tea/MedVLMBench
Results Specialist medical VLMs showed superior off-the-shelf performance on ID tasks, reflecting the benefits of targeted pretraining. However, lightweight fine-tuning enabled generalist VLMs to match or exceed specialist performance across most ID tasks. Fine-tuned generalist models also demonstrated stronger generalization on OOD datasets, whereas specialist models often exhibited degraded transfer performance.
Conclusion This benchmark provides the first systematic, paired comparison of generalist and specialist medical VLMs across diverse imaging domains. The findings indicate that efficiently adapted generalist VLMs can achieve competitive or superior performance relative to specialist models while offering improved flexibility and scalability for clinical AI development.
keywords:
Vision-language Models, Medical Image Analysis, Vision Question Answering, Diagnosis1 Introduction
@@ 3 Results
Together, these two views allow us to address all our research questions. The complete raw data underlying these figures are included in Appendix D, with Tables 5–7 reporting diagnostic results and Tables 5–5 reporting VQA results.
@@ 3.2 Comparison between generalist VLMs and the best-performing specialist medical VLMs
4 Discussion
This work provides a systematic head-to-head evaluation of generalistic and medical-specific VLMs across disease diagnosis and VQA, under explicitly defined ID and OOD settings and three evaluation regimes (off-the-shelf, lightweight adaptation on ID tasks, and lightweight adaptation on OOD tasks). Off-the-shelf, medical VLMs retain an advantage on ID datasets; however, after parameter-efficient fine-tuning, common VLMs meet or exceed specialist performance in most ID comparisons and show stronger transfer to OOD modalities, at a fraction of the training cost. This pattern is corroborated at the family level: for diagnosis and VQA, our RQ1–RQ3 framework isolates gaps between off-the-shelf models, then demonstrates how small-scale adaptation closes (or reverses) those gaps and improves OOD generalization. Representative results on FairVLMed10k show substantial gains after fine-tuning within Gemma and LLaVA families, illustrating that modest adaptation can equal or surpass medical counterparts across open-ended, closed-ended, and overall VQA metrics. Complementary analyses on OOD diagnosis further indicate that common models often generalize better than medical variants after tuning, consistent with the broader conclusion that efficient adaptation of strong generalist foundations is a cost-effective path to clinical imaging performance.
4.1 Limitations
Our ID/OOD definition is modality-based and does not capture finer distributional shifts (e.g., acquisition protocol, disease prevalence, institution-specific artifacts), which may also influence generalization. The benchmark, while broad, covers six diagnosis and four VQA datasets; expanding to additional modalities (e.g., ultrasound or nuclear medicine) and tasks with in-house datasets would provide a more exhaustive assessment. Our use of LLM-based judging (GPT Score) follows common practice for open-ended VQA but inherits known limitations of LLMs (e.g., occasional truncations or literal errors), which warrants cautious interpretation and continued exploration of reference-free automatic metrics and expert review. Finally, although our results consistently favor adapted generalist models, conclusions are conditioned on the datasets and hyperparameters considered here; different compute budgets, pretraining corpora, or training curricula for medical VLMs could shift relative performance.
4.2 Conclusions and Future Works
Across ID and OOD evaluations in diagnosis and VQA, efficiently fine-tuned general-purpose VLMs can match or surpass medical-specific VLMs, offering a scalable and cost-effective strategy without sacrificing performance or robustness. Future work will broaden the benchmark’s modality and task coverage, refine evaluation protocols (including stronger human and task-grounded assessments), and explore adaptation strategies that further enhance OOD generalization. In parallel, we envision data-centric extensions—curating more diverse multi-institutional cohorts and leveraging semi-automated labeling, active selection, and transfer learning—to reduce annotation cost while increasing realism and breadth, echoing successful dataset-scaling practices in related work.
@@ D.2 VQA
[H] Harvard-FairVLMed10k VQA performance - Qwen & Lingshu family. All values are shown as percentages. Each entry is the mean score, with the subscript showing the standard deviation, and parentheses indicating the confidence interval when available.
@@
[H] Harvard-FairVLMed10k VQA performance — Gemma family. All values are shown as percentages. Each entry is the mean score, with the subscript showing the standard deviation, and parentheses indicating the confidence interval when available.
@@
[H] Harvard-FairVLMed10k VQA performance — LLaVA family. All values are shown as percentages. Each entry is the mean score, with the subscript showing the standard deviation, and parentheses indicating the confidence interval when available.
@@
[H] Harvard-FairVLMed10k VQA performance — VILA family. All values are shown as percentages. Each entry is the mean score, with the subscript showing the standard deviation, and parentheses indicating the confidence interval when available.
@@
[H] Harvard-FairVLMed10k VQA performance — InternVL3 family. All values are shown as percentages. Each entry is the mean score, with the subscript showing the standard deviation, and parentheses indicating the confidence interval when available.
@@
[H] PathVQA VQA performance — Qwen & Lingshu family. All values are shown as percentages. Each entry is the mean score, with the subscript showing the standard deviation, and parentheses indicating the confidence interval when available.
@@
[H] PathVQA VQA performance — Gemma family. All values are shown as percentages. Each entry is the mean score, with the subscript showing the standard deviation, and parentheses indicating the confidence interval when available.
@@
[H] PathVQA VQA performance — LLaVA family. All values are shown as percentages. Each entry is the mean score, with the subscript showing the standard deviation, and parentheses indicating the confidence interval when available.
@@
[H] PathVQA VQA performance — VILA family. All values are shown as percentages. Each entry is the mean score, with the subscript showing the standard deviation, and parentheses indicating the confidence interval when available.
@@
[H] PathVQA VQA performance — InternVL3 family. All values are shown as percentages. Each entry is the mean score, with the subscript showing the standard deviation, and parentheses indicating the confidence interval when available.
@@
[H] SLAKE VQA performance — Qwen & Lingshu family. All values are shown as percentages. Each entry is the mean score, with the subscript showing the standard deviation, and parentheses indicating the confidence interval when available.
@@
[H] SLAKE VQA performance — Gemma family. All values are shown as percentages. Each entry is the mean score, with the subscript showing the standard deviation, and parentheses indicating the confidence interval when available.
@@
[H] SLAKE VQA performance — LLaVA family. All values are shown as percentages. Each entry is the mean score, with the subscript showing the standard deviation, and parentheses indicating the confidence interval when available.
@@
[H] SLAKE VQA performance — VILA family. All values are shown as percentages. Each entry is the mean score, with the subscript showing the standard deviation, and parentheses indicating the confidence interval when available.
@@
[H] SLAKE VQA performance — InternVL3 family. All values are shown as percentages. Each entry is the mean score, with the subscript showing the standard deviation, and parentheses indicating the confidence interval when available.
@@
[H] VQA-RAD VQA performance — Qwen & Lingshu family. All values are shown as percentages. Each entry is the mean score, with the subscript showing the standard deviation, and parentheses indicating the confidence interval when available.
@@
[H] VQA-RAD VQA performance — Gemma family. All values are shown as percentages. Each entry is the mean score, with the subscript showing the standard deviation, and parentheses indicating the confidence interval when available.
@@
[H] VQA-RAD VQA performance — LLaVA family. All values are shown as percentages. Each entry is the mean score, with the subscript showing the standard deviation, and parentheses indicating the confidence interval when available.
@@
[H] VQA-RAD VQA performance — VILA family. All values are shown as percentages. Each entry is the mean score, with the subscript showing the standard deviation, and parentheses indicating the confidence interval when available.
@@
[H] VQA-RAD VQA performance — InternVL3 family. All values are shown as percentages. Each entry is the mean score, with the subscript showing the standard deviation, and parentheses indicating the confidence interval when available.